OpenStudy (anonymous):

A tank of helium gas with a pressure of 240torr at 0 ∘C is heated to give a pressure of 1300torr

OpenStudy (cuanchi):

P1/T1=P2/T2 => T2= P2T1/P1 convert the temperature from C to K K=C+273.15 it is not necessary in this case convert the pressures to atm because it is going to be the same relationship (dividing both by the same number). calculate T2

OpenStudy (anonymous):

Thank you!